# Descent Methods for Vector Optimization Problems: A Majorization-Minimization Perspective

### v2

@@ 3.1 Majorization-minimization descent method for VOPs
In this section, we present a generic majorization-minimization scheme for minimizing a vector function in the sense of descent.
@@ 2       Choose a strongly $K$-convex surrogate function $G_{k}(\cdot)$ of $F(\cdot)-F(x^{k})$ near $x^{k}$
Algorithm 1 Generic_Majorization-minimization_descent_method_for_VOPs
@@
$\leq\left\langle{c^{*}_{k},F(x)-F(x^{k})}\right\rangle+\frac{1}{2}\left\lVert{x^{k}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle.$
This completes the proof. ∎
@@ 3.2.1 Global convergence
Firstly, we establish the global convergence result in the nonconvex cases, the following assumptions are required.
@@
    There exist two bases $C$ and $\tilde{C}$ such that
    $\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\leq\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\leq\max_{c^{*}\in\tilde{C}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle$
    holds for all $k\geq 0$. If $x^{k}\rightarrow x^{*}$, $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\rightarrow 0$, then $x^{*}$ is a $K$-stationary point to (VOP).
@@
The Assumption 3.1(i) is a standard condition for nonconvex cases. Moreover, the Assumption 3.1(ii) holds for many descent methods in vector (multiobjective) optimization, such as $K$-steepest descent method, Barzilai-Bowein descent method [7] and Newton method.
@@
The boundedness of $\{x^{k}\}$ indicates that $\{x^{k}\}$ has at least one accumulation point. Next, we prove that any accumulation point $x^{*}$ is a non-stationary point. By summing (4) from 00 to infinity, we have
@@
$-\infty<\max_{c^{*}\in C}\left\langle{c^{*},F^{*}-F(x^{0})}\right\rangle\leq\max_{c^{*}\in C}\left\langle{c^{*},\sum_{k=0}^{\infty}G_{k}(x^{k+1})}\right\rangle\leq\max_{c^{*}\in C_{k}}\left\langle{c^{*},\sum_{k=0}^{\infty}G_{k}(x^{k+1})}\right\rangle\leq\sum_{k=0}^{\infty}\max_{c^{*}\in\tilde{C}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle.$
This, together with the fact that $G_{k}(x^{k+1})\preceq_{K}G_{k}(x^{k})=0$, implies $\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\rightarrow 0$. For the accumulation point $x^{*}$, there exists an infinite index set $\mathcal{K}$ such that $x^{k}\stackrel{{\scriptstyle\mathcal{K}}}{{\longrightarrow}}x^{*}$. Therefore, by Assumption 3.1(ii), we conclude that $x^{*}$ is a $K$-stationary point. ∎
@@ 3.2.2 Strong convergence
In the following, we establish the strong convergence result of Algorithm 3.1 in $K$-convex cases.
@@
Suppose that Assumption 3.1 holds and $F(\cdot)$ is $K$-convex, let $\{x^{k}\}$ be the sequence generated by Algorithm 1 with $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $0\preceq_{K}\bm{\ell}\preceq_{K}\bm{\mu}$. Then, the following statements hold:
@@
    $u_{0}(x^{k})\leq\frac{\ell_{\max}R^{2}}{2k},~{}\forall k\geq 1$, where $\ell_{\max}:=\max_{c^{*}\in C_{k}}\left\langle{c^{*},\bm{\ell}}\right\rangle$, $R:=\{\left\lVert{x-y}\right\rVert:x,y\in\mathcal{L}_{F}(x^{0})\}$, and
    $u_{0}(x^{k}):=\max_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle$
@@
Furthermore, we use the fact that $0\preceq_{K}\bm{\ell}\preceq_{K}\bm{\mu}$ to get
@@
(ii) Since $0\preceq_{K}\bm{\ell}\preceq_{K}\bm{\mu}$, we use inequality (5) and the definition of $\ell_{\max}$ to obtain
$\left\langle{c^{*}_{k},F(x^{k+1})-F(x)}\right\rangle\leq\frac{1}{2}\left\lVert{x^{k}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle-\frac{1}{2}\left\lVert{x^{k+1}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle\leq\frac{\ell_{\max}}{2}\left(\left\lVert{x^{k}-x}\right\rVert^{2}-\left\lVert{x^{k+1}-x}\right\rVert^{2}\right).$
Taking the sum of the preceding inequality over 00 to $k-1$, we have
$\sum\limits_{s=0}^{k-1}\left\langle{c^{*}_{s},F(x^{s+1})-F(x)}\right\rangle\leq\frac{\ell_{\max}}{2}\left(\left\lVert{x^{0}-x}\right\rVert^{2}-\left\lVert{x^{k}-x}\right\rVert^{2}\right)\leq\frac{\ell_{\max}}{2}\left\lVert{x^{0}-x}\right\rVert^{2}.$
@@
$\sum\limits_{s=0}^{k-1}\left\langle{c^{*}_{s},F(x^{k})-F(x)}\right\rangle\leq\frac{\ell_{\max}}{2}\left\lVert{x^{0}-x}\right\rVert^{2}.$
Denote $\hat{c}^{*}_{k}:={\sum_{s=0}^{k-1}c^{*}_{s}}/{k}$, it follows by the convexity of $C_{k}$ that $\hat{c}^{*}_{k}\in C_{k}$. Therefore, we conclude that
$\left\langle{\hat{c}^{*}_{k},F(x^{k})-F(x)}\right\rangle\leq\frac{\ell_{\max}\left\lVert{x^{0}-x}\right\rVert^{2}}{2k}.$
Select $y^{k}\in\mathop{\arg\max}_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle$, it holds that
$u_{0}(x^{k})$ $=\max_{x\in\mathbb{R}^{n}}\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle=\min_{c^{*}\in C_{k}}\left\langle{c^{*},F(x^{k})-F(y^{k})}\right\rangle$
$\leq\left\langle{\hat{c}^{*}_{k},F(x^{k})-F(y^{k})}\right\rangle\leq\frac{\ell_{\max}\left\lVert{x^{0}-y^{k}}\right\rVert^{2}}{2k}.$
By the definition of $y^{k}$, we deduce that $y^{k}\in\{x:F(x)\preceq_{K}F(x^{k})\}\subset\mathcal{L}_{F}(x^{0})$, which implies $\left\lVert{x^{0}-z^{k}}\right\rVert\leq R$, the desired result follows. ∎
@@ 3.2.3 Linear convergence
Suppose that Assumption 3.1(ii) holds $F(\cdot)$ is strongly $K$-convex, let $\{x^{k}\}$ be the sequence generated by Algorithm 1 with $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $0\preceq_{K}\bm{\ell}\prec_{K}\bm{\mu}$. Then, the following statements hold:
@@
It seems that the convexity of $F(\cdot)$ plays no role in the proof of Theorems 3.2 and 3.3. However, it can indeed be shown that $F(\cdot)$ is necessarily $K$-convex if $\bm{\ell}=\bm{\mu}$ and strongly $K$-convex with $\bm{\mu}-\bm{\ell}$ if $\bm{\ell}\prec_{K}\bm{\mu}$. In the next section, we give some examples where such a condition holds.
@@
The linear convergence rate is related to $G_{k}(\cdot)$, which confirms that the rate of convergence can be improved by choosing a tighter surrogate.
4 Descent methods for VOPs without line search
In this section, we will revisit some well-known descent methods for VOPs and link them to the Algorithm 1. The main aim is to provide a unified theoretical analysis for all of them. We also develop some new algorithms with faster linear convergence.
@@ 4.1 $K$-steepest descent method for VOPs without line search
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{c^{*}\in C_{k}}\left\langle{c^{*},JF(x^{k})d}\right\rangle+\frac{1}{2}\|d\|^{2}.$ (6)
Select a vector $e\in{\rm int}(K)$, and denote $C_{e}=\{c^{*}\in K^{*}:\left\langle{c^{*},e}\right\rangle=1\}$. If we set $C_{k}=C_{e}$ in (6), then the $K$-steepest descent direction can be reformulated as the optimal solution of
@@
If $K=\mathbb{R}^{m}_{+}$, and $C_{e}=\Delta_{m}$, then $e=(1,...,1)^{T}$ and the subproblem (7) reduces to that of steepest descent method for MOPs [17].
@@
Algorithm 2 $K$-steepest_descent_method_without_line_search_for_VOPs
@@
By the definition of $L_{\max}$, we have $\bm{\ell}\preceq_{K}L_{\max}e$, it follows from the $K$-smoothness of $F(\cdot)$ that assertion (i) holds. Notice that $G_{k,Le}(\cdot)$ is strongly $K$-convex and $K$-smooth with $Le$, and $\bm{\mu}\preceq Le$, then we obtain assertion (ii) and (iii) by Lemma 2.3 (ii) and (iii), respectively. ∎
@@
Since $F(\cdot)$ is strongly $K$-convex, it follows that $G_{k,Le}(\cdot)\in\mathcal{S}_{Le-\bm{\mu},Le}(F,x^{k})$ and Assumption 3.1 holds in this case. By setting $C=C_{e}$, Theorem 3.3 (i) and (ii) reduce to the assertions (i) and (ii), respectively. ∎
@@
If $K=\mathbb{R}^{m}_{+}$ and $e=(1,...,1)^{T}$, then $C_{e}$ is a base of +$\mathbb{R}^{m}_{+}$, the convergence rate in Lemma 4.1 reduces to that of [44, Theorem 5.3] with $g(\cdot)=0$. It is worth noting that the rate of convergence is related to $e$, to the best of our knowledge, apart from $e=(1,...,1)^{T}$, it remains an open problem for the better choice of the parameter in MOPs.

### v3

@@ 3.1 Majorization-minimization descent method for VOPs
In this section, we present a unified majorization-minimization scheme for minimizing a vector-valued function in the sense of descent.
@@ 3   Choose a base $C_{k}$ of dual cone $K^{*}$
Algorithm 1 Unified majorization-minimization scheme for VOPs
Remark 3.1.
It is worth noting that we choose a variable base $C_{k}$ in each iteration, whreas an invariable base $C$ is used in existing descent methods for VOPs, see [22, 21, 20, 4, 33, 9].
@@
$\leq\left\langle{c^{*}_{k},F(x)-F(x^{k})}\right\rangle+\frac{1}{2}\left\lVert{x^{k}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle,$
where the equality follows by the definition of $H_{k}(\cdot)$ and the last inequality is due to the assertion (i). This completes the proof. ∎
@@ 3.2.1 Global convergence
Firstly, we establish the global convergence result in the nonconvex setting, the following assumptions are required.
@@
    There exist two compact bases $C_{L}$ and $\tilde{C}$ of $K^{*}$ such that
    $C_{k}\subset\tilde{C}$, and $\max_{c^{*}\in C_{k}}\left\langle{c^{*},y}\right\rangle\geq\max_{c^{*}\in C_{L}}\left\langle{c^{*},y}\right\rangle$
    hold for all $k\geq 0$ and $y\in-K$.
  • (iii)
    If $C$ is a compact base of $K^{*}$, $x^{k}\rightarrow x^{*}$, $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $\min_{y\in\mathbb{R}^{n}}\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(y)}\right\rangle\rightarrow 0$, then $x^{*}$ is a $K$-stationary point to (VOP).
@@
Assumption 3.1(i) is a standard condition for nonconvex cases. Moreover, Assumption 3.1(ii) requires the sequence $\{C_{k}\}$ to be uniformly bounded, which is a mild assumption. Assumption 3.1(iii) holds for $K$-steepest descent method [22]. Specifically, $\min_{y\in\mathbb{R}^{n}}\max_{c^{*}\in C}\left\langle{c^{*},G_{k}(y)}\right\rangle\rightarrow 0$ takes the form of $\alpha_{x_{k}}$ [22, Definition 3.2], where $\alpha_{x}$ is continuous [22, Lemma 3.3(3)] and $\alpha_{x}=0$ if and only if $x$ is a $K$-stationary point to (VOP) [22, Lemma 3.3(1)].
@@
The boundedness of $\{x^{k}\}$ indicates that $\{x^{k}\}$ has at least one accumulation point. Next, we prove that any accumulation point $x^{*}$ is a non-stationary point. By summing (7) from 0 to infinity, we have
@@
$\sum_{k=0}^{\infty}\max_{c^{*}\in{C}_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle$ $\geq\sum_{k=0}^{\infty}\max_{c^{*}\in{C}_{L}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle$
$\geq\max_{c^{*}\in C_{L}}\left\langle{c^{*},\sum_{k=0}^{\infty}G_{k}(x^{k+1})}\right\rangle$
$\geq\max_{c^{*}\in C_{L}}\left\langle{c^{*},F^{*}-F(x^{0})}\right\rangle$
$\geq-\infty,$
where the first inequality follows by Assumption 3.1(ii) and $G_{k}(x^{k+1})\preceq_{K}0$, the second inequality is due to the fact $\max_{x}{f_{1}(x)}+\max_{x}{f_{2}(x)}\geq\max_{x}\{f_{1}(x)+f_{2}(x)\}$. This, together with the fact that $G_{k}(x^{k+1})\preceq_{K}G_{k}(x^{k})=0$, implies $\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\rightarrow 0$. A direct calculation gives:
$0=\max_{c^{*}\in\tilde{C}}\left\langle{c^{*},G_{k}(x^{k})}\right\rangle\geq\min_{y\in\mathbb{R}^{n}}\max_{c^{*}\in\tilde{C}}\left\langle{c^{*},G_{k}(y)}\right\rangle\geq\min_{y\in\mathbb{R}^{n}}\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(y)}\right\rangle=\max_{c^{*}\in C_{k}}\left\langle{c^{*},G_{k}(x^{k+1})}\right\rangle\rightarrow 0,$
where the second inequality follows by $C_{k}\subset\tilde{C}$. Therefore, $\min_{y\in\mathbb{R}^{n}}\max_{c^{*}\in\tilde{C}}\left\langle{c^{*},G_{k}(y)}\right\rangle\rightarrow 0$. For the accumulation point $x^{*}$, there exists an infinite index set $\mathcal{K}$ such that $x^{k}\stackrel{{\scriptstyle\mathcal{K}}}{{\longrightarrow}}x^{*}$. By Assumption 3.1(iii), we conclude that $x^{*}$ is a $K$-stationary point. ∎
@@ 3.2.2 Strong convergence
In the following, we establish the strong convergence result of Algorithm 3.1 in $K$-convex setting.
@@
Suppose that Assumption 3.1 holds and $F(\cdot)$ is $K$-convex, let $\{x^{k}\}$ be the sequence generated by Algorithm 1 with $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $0\preceq_{K}\bm{\ell}=\bm{\mu}$. Then, the following statements hold:
@@
    $u_{0}(x^{k})\leq\frac{\ell_{\max}R^{2}}{2k},~\forall k\geq 1$, where $\ell_{\max}:=\max_{c^{*}\in\tilde{C}}\left\langle{c^{*},\bm{\ell}}\right\rangle$, $R:=\{\left\lVert{x-y}\right\rVert:x,y\in\mathcal{L}_{F}(x^{0})\}$, and
    $u_{0}(x^{k}):=\max_{x\in\mathbb{R}^{n}}\min_{c^{*}\in\tilde{C}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle$
@@
Furthermore, we use the fact that $0\preceq_{K}\bm{\ell}=\bm{\mu}$ to get
@@
(ii) Since $0\preceq_{K}\bm{\ell}=\bm{\mu}$, we use inequality (8) to obtain
$\left\langle{c^{*}_{k},F(x^{k+1})-F(x)}\right\rangle\leq\frac{1}{2}\left\lVert{x^{k}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle-\frac{1}{2}\left\lVert{x^{k+1}-x}\right\rVert^{2}\left\langle{c^{*}_{k},\bm{\mu}}\right\rangle=\frac{\left\langle{c^{*}_{k},\bm{\ell}}\right\rangle}{2}\left(\left\lVert{x^{k}-x}\right\rVert^{2}-\left\lVert{x^{k+1}-x}\right\rVert^{2}\right).$ (9)
Taking the sum of the preceding inequality over 0 to k1k-1, we have
$\sum\limits_{s=0}^{k-1}\left\langle{c^{*}_{s},F(x^{s+1})-F(x)}\right\rangle\leq\sum\limits_{s=0}^{k-1}\frac{\left\langle{c^{*}_{s},\bm{\ell}}\right\rangle}{2}\left(\left\lVert{x^{s}-x}\right\rVert^{2}-\left\lVert{x^{s+1}-x}\right\rVert^{2}\right).$
@@
$\sum\limits_{s=0}^{k-1}\left\langle{c^{*}_{s},F(x^{k})-F(x)}\right\rangle\leq\sum\limits_{s=0}^{k-1}\frac{\left\langle{c^{*}_{s},\bm{\ell}}\right\rangle}{2}\left(\left\lVert{x^{s}-x}\right\rVert^{2}-\left\lVert{x^{s+1}-x}\right\rVert^{2}\right).$
Denote $\hat{c}^{*}_{k}:={\sum_{s=0}^{k-1}c^{*}_{s}}/{k}$. It follows from the convexity of $\tilde{C}$ and the fact that $c^{*}_{s}\in\tilde{C}$ that $\hat{c}^{*}_{k}\in\tilde{C}$. Therefore, we conclude that
$\left\langle{\hat{c}^{*}_{k},F(x^{k})-F(x)}\right\rangle\leq\sum\limits_{s=0}^{k-1}\frac{\left\langle{c^{*}_{s},\bm{\ell}}\right\rangle}{2k}\left(\left\lVert{x^{s}-x}\right\rVert^{2}-\left\lVert{x^{s+1}-x}\right\rVert^{2}\right).$
Select $y^{k}\in\mathop{\arg\max}_{x\in\mathbb{R}^{n}}\min_{c^{*}\in\tilde{C}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle$, it holds that
$u_{0}(x^{k})$ $=\max_{x\in\mathbb{R}^{n}}\min_{c^{*}\in\tilde{C}}\left\langle{c^{*},F(x^{k})-F(x)}\right\rangle=\min_{c^{*}\in\tilde{C}}\left\langle{c^{*},F(x^{k})-F(y^{k})}\right\rangle$
$\leq\left\langle{\hat{c}^{*}_{k},F(x^{k})-F(y^{k})}\right\rangle\leq\sum\limits_{s=0}^{k-1}\frac{\left\langle{c^{*}_{s},\bm{\ell}}\right\rangle}{2k}\left(\left\lVert{x^{s}-y^{k}}\right\rVert^{2}-\left\lVert{x^{s+1}-y^{k}}\right\rVert^{2}\right).$
By the definition of $y^{k}$, we deduce that $y^{k}\in\{x:F(x)\preceq_{K}F(x^{k})\}\subset\mathcal{L}_{F}(x^{s})$ for all $s\leq k-1$. Substituting this relation into (9), we have $\left\lVert{x^{s}-y^{k}}\right\rVert^{2}-\left\lVert{x^{s+1}-y^{k}}\right\rVert^{2}\geq 0$ for all $s\leq k-1$. Therefore,
$u_{0}(x^{k})\leq\frac{\ell_{\max}}{2k}\sum\limits_{s=0}^{k-1}\left(\left\lVert{x^{s}-y^{k}}\right\rVert^{2}-\left\lVert{x^{s+1}-y^{k}}\right\rVert^{2}\right)\leq\frac{\ell_{\max}\left\lVert{x^{0}-y^{k}}\right\rVert^{2}}{2k}.$
Recall that $y^{k}\in\mathcal{L}_{F}(x^{0})$, the desired result follows. ∎
@@ 3.2.3 Linear convergence
Suppose that Assumption 3.1(ii) holds and $F(\cdot)$ is strongly $K$-convex, let $\{x^{k}\}$ be the sequence generated by Algorithm 1 with $G_{k}(\cdot)\in\mathcal{S}_{\bm{\ell},\bm{\mu}}(F,x^{k})$ and $0\preceq_{K}\bm{\ell}\prec_{K}\bm{\mu}$. Then, the following statements hold:
@@
[…]
@@
Note that $0\notin C_{k}$ and $\text{cone}(C_{k})=K^{*}$, we have
$\max\limits_{c^{*}\in C_{k}}\frac{\left\langle{c^{*},\bm{\ell}}\right\rangle}{\left\langle{c^{*},\bm{\mu}}\right\rangle}=\max\limits_{c^{*}\in K^{*}\setminus\{0\}}\frac{\left\langle{c^{*},\bm{\ell}}\right\rangle}{\left\langle{c^{*},\bm{\mu}}\right\rangle}.$
Therefore, the linear convergence rate is related to $\{G_{k}(\cdot)\}$, not $\{C_{k}\}$, which confirms that the rate of convergence can be improved by choosing a tighter surrogate.
4 First-order methods for VOPs with majorizing surrogate functions
It is worth noting that Remark 3.4 may suggest that the choice of the base in the subproblem is inessential, which could make the question (Q) raised in the introduction seem trivial. However, as will be demonstrated in this section, the selection of such a base not only influences the convergence rate but also determines the computational complexity of the subproblem. Both aspects are of central importance in the framework of majorization-minimization optimization.
In what follows, we first revisit the classical steepest descent method for VOPs (SDVO) [22] and establish its connection with Algorithm 1. To mitigate the slow convergence of SDVO, we then investigate, from a majorization-minimization perspective, how an appropriate choice of the base can effectively accelerate first-order methods for VOPs.
@@ 4.1 $K$-steepest descent method for VOPs without line search
$\min\limits_{d\in\mathbb{R}^{n}}\max\limits_{c^{*}\in C}\left\langle{c^{*},JF(x^{k})d}\right\rangle+\frac{1}{2}\|d\|^{2}.$ (10)
Select a vector $e\in{\rm int}(K)$, and denote $C_{e}=\{c^{*}\in K^{*}:\left\langle{c^{*},e}\right\rangle=1\}$. If we set $C=C_{e}$ in (10), then the $K$-steepest descent direction can be reformulated as the optimal solution of
@@
If $K=\mathbb{R}^{m}_{+}$, and $C_{e}=\Delta_{m}$, then $e=\mathbf{1}_{m}$ and the subproblem (11) reduces to that of steepest descent method for MOPs [15]. In what follows, we refer to subproblems of the forms (10) and (11) as seperate and coupled subproblems, respectively.
@@
Algorithm 2 $K$-steepest descent method for VOPs
@@
It is obvious that Algorithm 2 is a special case of Algorithm 1 with $C_{k}=C_{e}$ and $G_{k}(\cdot)=G_{k,Le}(\cdot)$. As described in Remark 3.4, the peformance of Algorithm 2 is mainly attributed to $G_{k,Le}(\cdot)$. The following results show that $G_{k,Le}(\cdot)$ is a majorizing surrogate function of $F(\cdot)-F(x^{k})$ near $x^{k}$.
@@
By the definition of $L_{\max}$, we have $\bm{\ell}\preceq_{K}L_{\max}e$, it follows from the $K$-smoothness of $F(\cdot)$ that assertion (i) holds. Notice that $G_{k,Le}(\cdot)$ is strongly $K$-convex and $K$-smooth with $Le$, and $\bm{\mu}\preceq_{K}Le$, then we obtain assertion (ii) and (iii) by Lemma 2.3 (ii) and (iii), respectively. ∎
Note that for a strongly $K$-convex objective function, $G_{k,Le}(\cdot)\in\mathcal{S}_{Le-\bm{\mu},Le}(F,x^{k})$ for all $L\geq L_{\max}$. We are now in a position to present the rate of linear convergence for SDVO.
@@
Since $F(\cdot)$ is strongly $K$-convex, it follows that $G_{k,Le}(\cdot)\in\mathcal{S}_{Le-\bm{\mu},Le}(F,x^{k})$ and Assumption 3.1 holds in this case. By setting $C_{k}=C_{e}$, Theorem 3.3 (i) and (ii) reduce to the assertions (i) and (ii), respectively. ∎
@@
If $K=\mathbb{R}^{m}_{+}$ and $e=\mathbf{1}_{m}$, then $C_{e}=\Delta_{m}$ is a base of $\mathbb{R}^{m}_{+}$, the convergence rate in Lemma 4.1 reduces to that of [45, Theorem 5.3] with $g(\cdot)=0$. Specifically, the linear convergence rate is worse than $\mathcal{O}((\sqrt{1-{\mu_{\min}}/{L_{\max}}})^{k})$ (setting $L=L_{\max}$), where $L_{\max}=\max_{i\in[m]}\{\bm{\ell}_{i}\}$ and $\mu_{\min}=\min_{i\in[m]}\{\bm{\mu}_{i}\}$. Therefore, even each of objective functions is well-conditioned ($\max_{i\in[m]}\{\bm{\ell}_{i}/\bm{\mu}_{i}\}$ is relative small), the linear convergence rate can be very slow due to objective imbalances ($L_{\max}/\mu_{\min}$ can be extremely large). It is worth noting that the rate of convergence is related to $C_{e}$, since in the seperate subproblem the surrogate function is inherently determined by $C_{e}$. To the best of our knowledge, apart from $\Delta_{m}$, it remains an open problem for the better choice of the base in MOPs.